In the Smart National Challenge, the probability of a person clearing the
certification, and appearing for the Equity Research certification is 0.5. There is
a record that given a candidate has appeared for an equity research
certification, the probability that he will clear it is 0.7. The chance of a person
getting in top 10 rank holders and appearing for the equity research certificate
is 0.3. What percentage of candidates, given they have appeared for an equity
research certification, are likely to get into top 10 rank holders?

Answer by stanbon(75887)

In the Smart National Challenge, the probability of a person clearing the
certification, and appearing for the Equity Research certification is 0.5. There is a record that, given a candidate has appeared for an equity research
certification, the probability that he will clear it is 0.7.
The chance of a person getting in top 10 rank holders and appearing for the equity research certificate s 0.3.
------------
What percentage of candidates, given they have appeared for an equity research certification, are likely to get into top 10 rank holders?
---
p(top 10 |appear) = P(top 10 AND appear)/P(appear) =
(0.5*0.7*0.3)/0.5 = 0.21
-----------------
Cheers,
Stan H.